Lea Lewis was an accountant in 1951 and earned $18,000 that year. Her son is an accountant too and he earned $240,000 this year. Suppose the price index was 18.1 in 1951 and 20.8 in the current year.

Refer to Scenario 24-1. In real terms, the income of Lea's son amounts to about what percentage of Lea's income?

a. 1160.3 percent

b. 13.3 percent

c. 8.6 percent

d. 1.1 percent
Answer #1

Real income of son = 240000 * 18.1 / 20.8 = 208846.15

% age of Lea income = 208846.15 / 18000 * 100 = 1160.256 % ~ 1160.3%

option a is correct answer